OpenStudy (anonymous):

The function f is defined by f(x) = x2 − 4x + 7 for x > 2. (i) Express f(x) in the form (x − a)2 + b and hence state the range of f. [3] (ii) Obtain an expression for f−1(x) and state the domain of f−1. [3] The function g is defined by g(x) = x − 2 for x > 2. The function h is such that f = hg and the domain of h is x > 0. (iii) Obtain an expression for h(x). [1]

OpenStudy (anonymous):

what is f?

OpenStudy (anonymous):

(i) (x-2)2+3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the range would be (3,infinity)

OpenStudy (anonymous):

(ii) f(x) = x2 − 4x + 6 domain all R

OpenStudy (anonymous):

(ii)\[f ^{-1}(x)=2+\sqrt{x-3}\] the domain \[[3,\infty]\]